# Can LLM-Simulated Practice and Feedback Upskill Human Counselors? A Randomized Study with 90+ Novice Counselors

### v1

@@ 5 Contributors
Ryan and Raj led the technical development efforts of CARE, the LLM-simulated training system. Ryan, Emma, and Diyi worked together to design the randomized experiment. Ryan, Ifdita, Juan Pablo, and Raj conducted the 90+ online video call sessions. Ryan was responsible for the automatic behavioral assessment of counseling skills, as well as the self-efficacy calibration analysis, while Emma and Diyi advised on analysis methods and interpretations of these quantitative results. Ifdita, Juan Pablo, and Ryan conducted the thematic analysis of participants qualitative, self-reflection data. All authors helped with drafting and editing the Article manuscript and figures, with a huge thanks for Raj, Diyi, and Emma for completing full editing passes over the Article.
6 Acknowledgments
This research was made possible with funding support from a Stanford Impact Labs Grant, Stanford HAI Seed Grant, Stanford Psychotherapy and Behavioral Sciences Department Innovator Award, and Stanford CURIS paid-internship program for undergraduate researchers.
@@ 6 Acknowledgments
The authors have a large community of researchers to thank for their help during all stages of this project. The CARE system would have not been possible without the significant code contributions from student researchers in SALT Lab, including Meijin Li, Cheng Chang, Ananjan Nandi, Alicja Chaszczewicz, and Alan Zhang. In addition, CARE’s design benefited from feedback from Bruce Arnow and others from the Stanford-PAU PsyD consortium. We thank Pannisy Zhao for assistance while conducting user studies. Statistical analyses and presentation of results were much improved due to help from Robert Kraut and Akhila Kovvuri. Thank you Yanzhe Zhang and Rose Wang for detailed discussions on finetuning the LLM classifiers. Thank you to the following individuals for providing feedback on early drafts and presentations of this work, with special thanks to Yutong Zhang, Will Held, Ella Li, Michael Ryan, Dora Zhao, Caleb Ziems, Matthew Jöerke, Joy He-Yueya, Omar Shaikh, and Kapil Garg.
@@ Appendix A Supplementary Materials
A.1 Prompts for CARE’s LLM-simulated Patients
The following patients were used during the randomized online lab study. Participants interacted with the same ordering of three AI patients in the pre-intervention assessment chat, intervention chat, and post-intervention chat.
These three patients were previously tested with domain-expert judges [15] and were rated an average of 6 out of 7 for realism. They were chosen to balance (1) diversity across different ages, genders, and presenting counseling concern; and (2) consistency in challenging behaviors. In particular, each patient’s prompt defined similar behavioral principles to either not suggest solutions on one’s own, or to show a degree of skepticism or reluctance to a therapist’s advice or solutions. Thus, these AI patients were designed to challenge the novice counselors in similar ways, despite the diversity in scenarios. We highlight in blue these principles that resemble this resistance to arriving at or accepting solutions.
A.1.1 AI Patient for Pre-intervention Assessment Chat
@@ 7. Dont be so self-aware or good at recognizing your own problems@
A.1.2 AI Patient for Intervention Chat
@@
The member is a 35-year-old male, cisgender, heterosexual
veteran who has recently presented to treatment to address
his substance use issues and legal issues. He is court-
mandated to therapy. He had a severe psychotic break in
the context of marijuana and psilocybin about 6 months ago,
where he experienced paranoid delusions and hallucinations.
He is now stable and is not experiencing any psychotic
symptoms. In therapy, he is hoping to work on his estranged
relationship with his parents, who are currently caring for
the members two young children. The member is adamantly
focused on being reunited with his parents. In therapy, the
member is unable to accept that he had a recent episode of
psychosis and is very resistant to anything that resembles
criticism. He does not view himself as having any issue and
believes that all of the problems in his life are because
of other people. His demeanor is hostile and somewhat
aggressive, and he is quick to shut down any conversation
that might identify his own areas of development. He
struggles to feel emotions beyond anger and frustration.
@@ 4. Dont be so self-aware or good at recognizing your own problems
A.1.3 AI Patient for Post-intervention Assessment Chat
@@ 8. Dont be so self-aware or good at recognizing your own problems@
Figure 6: Overview of the core counseling skills introduced during the 5-minute static tutorial. Tutorial included 8 core counseling skills—such as reflections, empathy, validation, and suggestions—with definitions, usage tips, and example responses. This tutorial was provided to participants prior to engaging in simulated counseling practice.
@@
Table 3: To study the Dunning-Kruger effect and the change in discrepancy between perceived and actual ability, we map specific self-efficacy factors to corresponding NLP-based behavioral assessments.
@@
Table 4: Testing for Dunning-Kruger effects for Exploration and Action Skills using the classic quartile ANOVA analysis. Notes: $Pre(All)$ denotes all 94 participants’ assessments for the pre-chat, while $Post(All)$ is the same measured for the post-chat. $\eta^{2}_{g}$=generalized eta squared. * indicates significance after Bonferroni correction
@@
Timepoint Quartile $t$ $df$ $M_{diff}$ $95\%$ BCa CI $p$ $d$
Pre 1 -7.46 22.00 -50.38 [-63.34; -37.26] <0.001<0.001*< 0.001 ∗ -1.56
Pre 2 -6.23 24.00 -36.59 [-47.61; -25.42] <0.001<0.001*< 0.001 ∗ -1.25
Pre 3 -9.29 21.00 -43.57 [-52.35; -34.58] <0.001<0.001*< 0.001 ∗ -1.98
Pre 4 -5.94 23.00 -34.97 [-45.84; -23.40] <0.001<0.001*< 0.001 ∗ -1.21
Post 1 -7.86 23.00 -47.17 [-58.56; -35.32] <0.001<0.001*< 0.001 ∗ -1.60
Post 2 -7.63 21.00 -51.16 [-63.63; -37.96] <0.001<0.001*< 0.001 ∗ -1.63
Post 3 -8.02 24.00 -44.53 [-55.08; -34.11] <0.001<0.001*< 0.001 ∗ -1.60
Post 4 -8.66 22.00 -47.08 [-58.13; -37.39] <0.001<0.001*< 0.001 ∗ -1.81
@@
Timepoint Quartile $t$ $df$ $M_{diff}$ $95\%$ BCa CI $p$ $d$
Pre 1 -10.24 21.00 -60.05 [-70.80; -48.12] <0.001<0.001*< 0.001 ∗ -2.18
Pre 2 -7.37 26.00 -42.34 [-52.88; -31.43] <0.001<0.001*< 0.001 ∗ -1.42
Pre 3 -5.51 15.00 -35.00 [-47.87; -23.58] <0.001<0.001*< 0.001 ∗ -1.38
Pre 4 -6.54 27.00 -33.00 [-42.70; -23.38] <0.001<0.001*< 0.001 ∗ -1.24
Post 1 -10.50 23.00 -58.48 [-69.50; -47.99] <0.001<0.001*< 0.001 ∗ -2.14
Post 2 -6.45 19.00 -43.27 [-56.17; -31.20] <0.001<0.001*< 0.001 ∗ -1.44
Post 3 -7.09 30.00 -35.28 [-44.63; -25.88] <0.001<0.001*< 0.001 ∗ -1.27
Post 4 -9.15 18.00 -53.13 [-64.08; -42.01] <0.001<0.001*< 0.001 ∗ -2.10
Table 5: Pairwise Comparisons of Self-Efficacy and Performance Percentiles by Quartile and Timepoint. Note: Bootstrapped paired t-tests comparing self-efficacy and performance percentiles across quartiles. * indicates significance after Bonferroni correction.
@@
Timepoint Group Quartile $t$ $df$ $M_{diff}$ $95\%$ BCa CI $p$ $d$
@@
Pre P 4 -4.16 16.00 -29.97 [-44.70; -16.70] <0.001<0.001*< 0.001 ∗ -1.01
Pre P+F 1 -7.43 10.00 -60.55 [-75.96; -44.89] <0.001<0.001*< 0.001 ∗ -2.24
Pre P+F 2 -4.35 14.00 -33.65 [-48.45; -19.85] <0.001<0.001*< 0.001 ∗ -1.12
Pre P+F 3 -7.57 13.00 -37.18 [-46.58; -27.60] <0.001<0.001*< 0.001 ∗ -2.02
Pre P+F 4 -5.11 6.00 -47.11 [-64.91; -32.75] <0.001<0.001*< 0.001 ∗ -1.93
Post P 1 -4.75 12.00 -38.33 [-53.51; -22.57] <0.001<0.001*< 0.001 ∗ -1.32
Post P 2 -7.19 13.00 -58.16 [-72.71; -41.81] 0.0010.001*0.001 ∗ -1.92
Post P 3 -6.06 12.00 -43.11 [-56.36; -29.54] <0.001<0.001*< 0.001 ∗ -1.68
@@
Post P+F 1 -6.99 10.00 -57.63 [-73.04; -42.12] 0.0030.003*0.003 ∗ -2.11
Post P+F 2 -3.48 7.00 -38.91 [-59.70; -20.19] 0.0040.004*0.004 ∗ -1.23
Post P+F 3 -5.15 11.00 -46.06 [-62.73; -29.62] 0.0010.001*0.001 ∗ -1.49
Post P+F 4 -7.39 15.00 -42.86 [-53.72; -31.96] <0.001<0.001*< 0.001 ∗ -1.85
@@
Timepoint Group Quartile $t$ $df$ $M_{diff}$ $95\%$ BCa CI $p$ $d$
Pre P 1 -7.49 7.00 -65.87 [-81.33; -49.73] 0.0040.004*0.004 ∗ -2.65
Pre P 2 -5.05 15.00 -43.23 [-59.33; -27.46] <0.001<0.001*< 0.001 ∗ -1.26
Pre P 3 -3.67 9.00 -34.31 [-52.11; -17.97] 0.0050.005*0.005 ∗ -1.16
@@
Pre P+F 4 -3.94 14.00 -26.57 [-40.34; -14.92] 0.0010.001*0.001 ∗ -1.02
Post P 1 -6.90 11.00 -57.56 [-72.21; -41.60] 0.0020.002*0.002 ∗ -1.99
Post P 2 -5.79 9.00 -54.08 [-71.82; -36.12] 0.0030.003*0.003 ∗ -1.83
Post P 3 -3.81 14.00 -29.72 [-44.60; -15.85] 0.0010.001*0.001 ∗ -0.98
Post P 4 -7.62 9.00 -59.76 [-73.85; -45.51] <0.001<0.001*< 0.001 ∗ -2.41
Post P+F 1 -7.67 11.00 -59.39 [-72.88; -44.61] <0.001<0.001*< 0.001 ∗ -2.21
Post P+F 2 -3.70 9.00 -32.46 [-48.86; -17.08] 0.0010.001*0.001 ∗ -1.17
Post P+F 3 -6.47 15.00 -40.48 [-52.48; -28.69] <0.001<0.001*< 0.001 ∗ -1.62
Post P+F 4 -5.46 8.00 -45.76 [-62.25; -31.00] 0.0020.002*0.002 ∗ -1.82
Table 6: Pairwise Comparisons of Self-Efficacy and Performance Percentiles by Group and Quartile, measured for the pre-assessment chat and post-assessment chat. Note: Bootstrapped paired t-tests comparing self-efficacy and performance percentiles across quartiles at pre-test and post-test. * indicates significance after Bonferroni correction.
@@
Table 7: Qualitative Coding of Open-Ended Reflections of P + F Group Participants
@@
Table 8: Qualitative Coding of Open-Ended Reflections of P Group Participants

### v2

@@ 9. Contributors
Ryan and Raj led the technical development efforts of CARE, the LLM-simulated training system. Ryan, Emma, and Diyi worked together to design the randomized experiment. Ryan, Ifdita, Juan Pablo, and Raj conducted the 90+ online video call sessions. Ryan was responsible for the automatic behavioral assessment of counseling skills, as well as the self-efficacy calibration analysis, while Emma and Diyi advised on analysis methods and interpretations of these quantitative results. Ifdita, Juan Pablo, and Ryan conducted the thematic analysis of participants qualitative, self-reflection data. Ryan conducted the in-depth analysis of participant perceptions of receiving LLM feedback and practicing with simulated patients. All authors helped with drafting and editing the Article manuscript, references, and figures.
Acknowledgements.
@@
This research was made possible with funding support from a Stanford Impact Labs Stage 1 Award, Stanford HAI Seed Grant, Stanford Psychotherapy and Behavioral Sciences Department Innovator Award, and Stanford CURIS paid-internship program for undergraduate researchers. The authors have a large community of researchers to thank for their help during all stages of this project. The CARE system would have not been possible without the significant code contributions from student researchers in SALT Lab, including Meijin Li, Cheng Chang, Ananjan Nandi, Alicja Chaszczewicz, and Alan Zhang. In addition, CARE’s design benefited from feedback from Bruce Arnow and others from the Stanford-PAU PsyD consortium. We thank Pannisy Zhao for assistance while conducting user studies. Statistical analyses and presentation of results were much improved due to help from Robert Kraut and Akhila Kovvuri. Thank you Yanzhe Zhang and Rose Wang for detailed discussions on finetuning the LLM classifiers. Thank you to the following individuals for providing feedback on early drafts and presentations of this work, with special thanks to Yutong Zhang, Will Held, Ella Li, Michael Ryan, Dora Zhao, Caleb Ziems, Matthew Jöerke, Joy He-Yueya, Omar Shaikh, and Kapil Garg.
@@ Appendix A Supplementary Materials
A.1. Prompts for CARE’s LLM-simulated Patients
Tables 45, and 6 contain the profile of the three AI patients used in the randomized online lab study; participants interacted with them in the same order across the pre-intervention assessment chat, intervention chat, and post-intervention chat. Previously validated by domain experts (Louie et al., 2024), they averaged 6 out of 7 for realism. They were chosen to balance diversity (age, gender, and presenting concern) with consistency in challenging behaviors: each prompt enforces similar principles—such as reluctance to offer or accept solutions—so novice counselors encounter comparable resistance despite varied scenarios. We highlight in blue these principles that resemble this resistance to arriving at or accepting solutions.
Table 4. AI Patient for Pre-intervention Assessment Chat
@@ 7. Dont be so self-aware or good at recognizing your own problems@
Table 5. AI Patient for Intervention Chat
@@
The member is a 35-year-old male, cisgender, heterosexual veteran who has recently presented to treatment to address his substance use issues and legal issues. He is court-mandated to therapy. He had a severe psychotic break in the context of marijuana and psilocybin about 6 months ago, where he experienced paranoid delusions and hallucinations. He is now stable and is not experiencing any psychotic symptoms. In therapy, he is hoping to work on his estranged relationship with his parents, who are currently caring for the members two young children. The member is adamantly focused on being reunited with his parents. In therapy, the member is unable to accept that he had a recent episode of psychosis and is very resistant to anything that resembles criticism. He does not view himself as having any issue and believes that all of the problems in his life are because of other people. His demeanor is hostile and somewhat aggressive, and he is quick to shut down any conversation that might identify his own areas of development. He struggles to feel emotions beyond anger and frustration.
@@ 4. Dont be so self-aware or good at recognizing your own problems
Table 6. AI Patient for Post-intervention Assessment Chat
@@ 8. Dont be so self-aware or good at recognizing your own problems@
Figure 6. Overview of the core counseling skills introduced during the 5-minute static tutorial. The tutorial included 8 core counseling skills, such as reflections, empathy, validation, and suggestions, with definitions, usage tips, and example responses. This tutorial was provided to participants prior to engaging in simulated counseling practice.
The figure shows an overview of the core counseling skills introduced during the 5-minute static tutorial. The tutorial included eight core counseling skills, such as reflections, empathy, validation, and suggestions, with definitions, usage tips, and example responses. This tutorial was provided to participants before engaging in simulated counseling practice.
A.2. Classifier Selection Criteria
From the initial set of 16 binary classifiers, we down-selected to four final classifiers (strong uses of Empathy, Reflections, and Questions, plus areas needing improvement for Suggestions) based on the following criteria:
  1. (1)
    Performance metrics: Focusing on classifiers with the strongest performance metrics. Our selected classifiers achieved F1 scores ranging from 0.507 to 0.775, representing the highest-performing subset from our validation results.
  2. (2)
    Statistical considerations: The need to limit the number of statistical comparisons to avoid diluting significance across too many tests. With four classifiers, we conducted 12 planned t-tests (4 classifiers × 3 analyses each: within-group changes for P, within-group changes for P+F, and between-group differences).
  3. (3)
    Theoretical relevance: Selecting skills that represent core competencies in client-centered frameworks and are frequently used in counseling sessions. The chosen skills span both exploration stage (Empathy, Reflections, Questions) and action stage (Suggestions) of Hill’s Helping Skills framework.
Excluded classifiers: Self-disclosure was excluded from our analysis due to its infrequency in our dataset (appearing in fewer than 5% of utterances). Validation, though conceptually related to Empathy and mentioned frequently in qualitative data, showed more limited classifier performance (F1=0.556 for strengths) and was therefore reserved for secondary analyses. Session Management and Professionalism were excluded from fine-tuning entirely due to infrequent occurrence in the training data.
A.3. Changes in Self-Efficacy and Calibration with Behavioral Performance
See Tables 789, and 10.
@@ A.3. Changes in Self-Efficacy and Calibration with Behavioral Performance
Table 7. To study the Dunning-Kruger effect and the change in discrepancy between perceived and actual ability, we map specific self-efficacy factors to corresponding NLP-based behavioral assessments.
@@
Table 8. Testing for Dunning-Kruger effects for Exploration and Action Skills using the classic quartile ANOVA analysis. Notes: $Pre(All)$ denotes all 94 participants’ assessments for the pre-chat, while $Post(All)$ is the same measured for the post-chat. $\eta^{2}_{g}$=generalized eta squared. * indicates significance after Bonferroni correction
@@
Timepoint Quartile $t$ dfdf $M_{diff}$ $95\%$ BCa CI $p$ $d$
Pre 1 -7.46 22.00 -50.38 [-63.34; -37.26] <0.001<0.001* -1.56
Pre 2 -6.23 24.00 -36.59 [-47.61; -25.42] <0.001<0.001* -1.25
Pre 3 -9.29 21.00 -43.57 [-52.35; -34.58] <0.001<0.001* -1.98
Pre 4 -5.94 23.00 -34.97 [-45.84; -23.40] <0.001<0.001* -1.21
Post 1 -7.86 23.00 -47.17 [-58.56; -35.32] <0.001<0.001* -1.60
Post 2 -7.63 21.00 -51.16 [-63.63; -37.96] <0.001<0.001* -1.63
Post 3 -8.02 24.00 -44.53 [-55.08; -34.11] <0.001<0.001* -1.60
Post 4 -8.66 22.00 -47.08 [-58.13; -37.39] <0.001<0.001* -1.81
@@
Timepoint Quartile $t$ dfdf $M_{diff}$ $95\%$ BCa CI $p$ $d$
Pre 1 -10.24 21.00 -60.05 [-70.80; -48.12] <0.001<0.001* -2.18
Pre 2 -7.37 26.00 -42.34 [-52.88; -31.43] <0.001<0.001* -1.42
Pre 3 -5.51 15.00 -35.00 [-47.87; -23.58] <0.001<0.001* -1.38
Pre 4 -6.54 27.00 -33.00 [-42.70; -23.38] <0.001<0.001* -1.24
Post 1 -10.50 23.00 -58.48 [-69.50; -47.99] <0.001<0.001* -2.14
Post 2 -6.45 19.00 -43.27 [-56.17; -31.20] <0.001<0.001* -1.44
Post 3 -7.09 30.00 -35.28 [-44.63; -25.88] <0.001<0.001* -1.27
Post 4 -9.15 18.00 -53.13 [-64.08; -42.01] <0.001<0.001* -2.10
Table 9. Pairwise Comparisons of Self-Efficacy and Performance Percentiles by Quartile and Timepoint. Note: Bootstrapped paired t-tests comparing self-efficacy and performance percentiles across quartiles. * indicates significance after Bonferroni correction.
@@
Timepoint Group Quartile $t$ dfdf $M_{diff}$ $95\%$ BCa CI $p$ $d$
@@
Pre P 4 -4.16 16.00 -29.97 [-44.70; -16.70] <0.001<0.001* -1.01
Pre P+F 1 -7.43 10.00 -60.55 [-75.96; -44.89] <0.001<0.001* -2.24
Pre P+F 2 -4.35 14.00 -33.65 [-48.45; -19.85] <0.001<0.001* -1.12
Pre P+F 3 -7.57 13.00 -37.18 [-46.58; -27.60] <0.001<0.001* -2.02
Pre P+F 4 -5.11 6.00 -47.11 [-64.91; -32.75] <0.001<0.001* -1.93
Post P 1 -4.75 12.00 -38.33 [-53.51; -22.57] <0.001<0.001* -1.32
Post P 2 -7.19 13.00 -58.16 [-72.71; -41.81] 0.0010.001* -1.92
Post P 3 -6.06 12.00 -43.11 [-56.36; -29.54] <0.001<0.001* -1.68
@@
Post P+F 1 -6.99 10.00 -57.63 [-73.04; -42.12] 0.0030.003* -2.11
Post P+F 2 -3.48 7.00 -38.91 [-59.70; -20.19] 0.0040.004* -1.23
Post P+F 3 -5.15 11.00 -46.06 [-62.73; -29.62] 0.0010.001* -1.49
Post P+F 4 -7.39 15.00 -42.86 [-53.72; -31.96] <0.001<0.001* -1.85
@@
Timepoint Group Quartile $t$ dfdf $M_{diff}$ $95\%$ BCa CI $p$ $d$
Pre P 1 -7.49 7.00 -65.87 [-81.33; -49.73] 0.0040.004* -2.65
Pre P 2 -5.05 15.00 -43.23 [-59.33; -27.46] <0.001<0.001* -1.26
Pre P 3 -3.67 9.00 -34.31 [-52.11; -17.97] 0.0050.005* -1.16
@@
Pre P+F 4 -3.94 14.00 -26.57 [-40.34; -14.92] 0.0010.001* -1.02
Post P 1 -6.90 11.00 -57.56 [-72.21; -41.60] 0.0020.002* -1.99
Post P 2 -5.79 9.00 -54.08 [-71.82; -36.12] 0.0030.003* -1.83
Post P 3 -3.81 14.00 -29.72 [-44.60; -15.85] 0.0010.001* -0.98
Post P 4 -7.62 9.00 -59.76 [-73.85; -45.51] <0.001<0.001* -2.41
Post P+F 1 -7.67 11.00 -59.39 [-72.88; -44.61] <0.001<0.001* -2.21
Post P+F 2 -3.70 9.00 -32.46 [-48.86; -17.08] 0.0010.001* -1.17
Post P+F 3 -6.47 15.00 -40.48 [-52.48; -28.69] <0.001<0.001* -1.62
Post P+F 4 -5.46 8.00 -45.76 [-62.25; -31.00] 0.0020.002* -1.82
Table 10. Pairwise Comparisons of Self-Efficacy and Performance Percentiles by Group and Quartile, measured for the pre-assessment chat and post-assessment chat. Note: Bootstrapped paired t-tests comparing self-efficacy and performance percentiles across quartiles at pre-test and post-test. * indicates significance after Bonferroni correction.
@@
Table 11. Qualitative Coding of Open-Ended Reflections of P + F Group Participants
@@
Table 12. Qualitative Coding of Open-Ended Reflections of P Group Participants